OpenStudy (anonymous):

If f(x) = 2x(to the power of 2) + 2, find f(5).

OpenStudy (anonymous):

Plug in 5 into every x in you function so \[f(5)=(2*5)^2\] or \[f(5)=2*5^2\] since I am not sure what part of the function is raised to the power of 2.